### v3

# DurFlex-EVC: Duration-Flexible Emotional Voice Conversion with Parallel Generation

Abstract
Emotional voice conversion involves modifying the pitch, spectral envelope, and other acoustic characteristics of speech to match a desired emotional state while maintaining the speaker’s identity. Recent advances in EVC involve simultaneously modeling pitch and duration by exploiting the potential of sequence-to-sequence models. In this study, we focus on parallel speech generation to increase the reliability and efficiency of conversion. We introduce a duration-flexible EVC (DurFlex-EVC) that integrates a style autoencoder and a unit aligner. The previous variable-duration parallel generation model required text-to-speech alignment. We consider self-supervised model representation and discrete speech units to be the core of our parallel generation. The style autoencoder promotes content style disentanglement by separating the source style of the input features and applying them with the target style. The unit aligner encodes unit-level features by modeling emotional context. Furthermore, we enhance the style of the features with a hierarchical stylize encoder and generate high-quality Mel-spectrograms with a diffusion-based generator. The effectiveness of the approach has been validated through subjective and objective evaluations and has been demonstrated to be superior to baseline models.
Index Terms:
emotional voice conversion, self-supervised representation, style disentanglement, duration control
I Introduction
Emotional voice conversion (EVC) involves modifying various acoustic characteristics of a voice, such as pitch and spectral envelope, to match a desired emotional state while preserving the speaker’s identity. EVC has gained prominence, particularly in the realm of voice-interactive technologies such as virtual assistants and internet of things (IoT) devices, improving the human-like and emotionally resonant aspects of digital interactions [1, 2, 3, 4].
In the context of EVC, a crucial objective is to preserve the speaker identity and content of the original speech while modifying only those speech attributes that convey emotion [5, 6]. This necessitates an adjustment of the prosody to align with the intended emotion. Prosodic elements, including intonation, rhythm, and energy, play a critical role in both conveying and recognizing emotions in speech. Although the concept of controlling prosody for emotional conversion is intuitively appealing [7], the process of refining each prosodic component presents a significant challenge.
The field of EVC has been revolutionized by advances in deep learning [8, 9]. Some studies employed Gaussian mixture models [10] to convert spectral and prosodic features to produce more expressive voices. Subsequent developments led to autoencoder-based methods [11, 12, 13], enabling learning in non-parallel data-driven EVC. To convert non-parallel emotional speech without changing the speaker’s identity and linguistic content, some VAE-based methods [14] have been proposed. GAN-based approaches [15], using frameworks such as Cycle-GAN [16], StarGAN [17], and VAE-GAN [18], represent further advances. However, these methods often overlook the importance of rhythm when expressing emotion because they support emotional conversion with a fixed length.
Sequence-to-sequence (Seq2Seq)-based models, capable of implicitly modeling duration, have emerged as a significant development [19, 20]. These models often adopt specific strategies, such as a two-stage learning strategy integrating a text-to-speech (TTS) model, to improve learning stability [21, 22]. Although seq2seq models can generate varying durations, they face typical autoregressive model challenges, such as long-term dependency and repetition issues. This necessitates a parallel generation approach for efficiency and reliability. In this study, we explored parallel generation methods with flexible durations. To enable parallel generation, the duration of the content needs to be explicitly modeled. Some voice conversion models [23] have leveraged phoneme duration from text-to-speech models. Obtaining phoneme durations requires additional effort, such as using the encoder-decoder attention of a pre-trained autoregressive TTS or using an external forced alignment tool.
Recently, the exploration of discrete speech units through self-supervised learning representations has shown promise in addressing parallel generation challenges in speech processing. This technique encodes speech into discrete units, facilitating frame-level duration extraction for each unit. Certain studies [24, 25] have investigated leveraging these properties. For instance, [24, 25] proposed an approach similar to spoken language translation for speech emotion conversion, allowing parallel audio generation by predicting the duration of each unit. However, this approach does not fully achieve parallel generation due to its reliance on autoregressive models for emotion translation.
In this paper, we propose a duration-flexible EVC (DurFlex-EVC) that supports parallel generation. We consider discrete speech units as the key to parallel generation and duration flexibility. Although the units have content information, the acoustic characteristics are not sufficient, and thus we use other speech representations as input instead. Flexible duration is achieved by guiding the model to predict units, deduplicating features into unique units, and scaling to frame level by predicting unit durations corresponding to target emotions. Our main contributions are as follows.
  • We achieved flexible duration modeling using discrete speech units and introduced unit aligners to model emotional stylistic context.
  • We designed a style autoencoder to disentangle the content and emotional style of input features. The style autoencoder de-stylizes source emotion from input features and stylizes target emotion for the desired emotion.
  • We introduced a hierarchical stylize encoder that adapts styles at the unit-level and frame-level.
  • We adopted a diffusion-based generator to produce high-quality speech.
  • We propose emotion embedding cosine similarity (EECS) as a method to objectively evaluate the emotional expressiveness of generated speech.
  • We conduct comprehensive evaluations to demonstrate the effectiveness of each component and their contributions to the overall performance of the model.
II Background
II-A Exploring Self-Supervised Learning in Speech
Self-supervised learning (SSL) is a machine learning paradigm in which models are trained on their own datasets to create meaningful representations. This method is particularly beneficial in speech processing, where data labeling can be both time-consuming and costly. The wav2vec 2.0 model [26] used contrastive learning to validate SSL representations. The vq-wav2vec model [27] introduced a technique to learn discrete audio representations through self-supervised context prediction and quantization. XLS-R [28] is an extensive cross-lingual speech representation model based on wav2vec 2.0. Hidden-unit BERT (HuBERT) [29] employs a masked prediction approach similar to BERT [30] for learning representations. ContentVec [31] improves speaker disentanglement within the HuBERT framework. Efforts have also been made to create representations suitable for universal downstream tasks [32].
Recently, SSL representations have been extensively applied to various downstream tasks, such as automatic speech recognition [26], voice conversion [33], speaker verification [34], speech synthesis [35], speech emotion recognition [36], and speech enhancement [37].
II-B Discrete Units in Speech Processing
In the realm of audio and speech, discrete unit representation has been proposed for diverse tasks. SoundStream [38] introduced a neural audio codec employing a residual vector quantizer (RVQ), while EnCodec [39] focuses on high-fidelity audio compression and lightweighting through similar methods. UniAudio [40] emerged as a general-purpose audio generation model. These neural codec-based methods, aimed primarily at audio compression and restoration, feature large codebooks and relatively small dimensions.
In contrast, certain methods emphasize the compression of speech into semantic units. A method was proposed to decompose and reconstruct speech into discrete units of pitch and speaker identity [41]. Based on this, soft speech units were suggested [24] for enhanced content capture, thereby improving the naturalness and intelligibility. Furthermore, speech emotion conversion was explored as a language translation task [42], using discrete representations of phonetic content, prosody, speaker, and emotion in conjunction with a neural vocoder for waveform generation. UnitSpeech [25] demonstrated proficiency in personalized TTS and voice conversion, fine-tuning a diffusion-based TTS model with minimal data using self-supervised units, eliminating the need for retraining for each task. Here, semantic speech units serve as content in speech decomposition.
II-C Parallel Speech Generation
Non-autoregressive speech synthesis models generate speech frames in parallel, significantly reducing inference time compared to autoregressive methods. In a recent speech synthesis study, parallel generation methods outperformed and were more reliable than autoregressive methods for text-to-speech, voice conversion, and vocoder. For parallel generation, TTS requires an alignment between text and speech, which can be extracted from pre-trained autoregressive teacher models [43], using an external aligner [44] such as the Montreal Forced Aligner (MFA) [45], or using the monotonic alignment search algorithm (MAS) [46]. For voice conversion, many approaches [47, 48] employ parallel generation, but some models adopt autoregressive [49] due to the impossibility of converting duration. Some works [50, 23] have proposed models with variable duration and parallel generation; these models use text-speech alignment as in TTS. For vocoders, studies have been proposed to upsample the input features and then generate waveforms through various generation models such as GAN [51], normalizing flow [52], or generative diffusion model[53]. Inspired by the success of parallel generation models, we aimed to adopt a parallel generation framework for emotional voice conversion.
II-D Duration Modeling in Speech Processing
Duration modeling is a critical aspect of speech synthesis, particularly in TTS, where mismatches between character length and signals can occur. Early TTS methods addressed this through autoregressive models with implicit duration modeling, generating one frame at a time [54, 55]. FastSpeech [43] leveraged the encoder-decoder attention alignment of an autoregressive teacher model to model phoneme duration, facilitating parallel generation. FastSpeech 2 [44] introduced a method for extracting phoneme duration from forced external alignment. Glow-TTS [46] developed a technique for identifying the most similar monotonic alignment between text and latent using a monotonic alignment search. Duration modeling advances have also been applied to voice conversion, with seq2seq models handling duration changes [56, 49]. The DCVC model [23] utilized a phoneme-based information bottleneck for style transfer and speech speed control in voice conversion. Discrete speech units have been employed [41, 25] to model duration through consecutive unit counts. In EVC, a trend towards seq2seq structures for handling duration changes has been observed [22, 57]. However, despite [42] proposing EVC using units for parallel generation, it still relies on the seq2seq model for the unit translation process. We leveraged units to model duration without alignment to text-speech to facilitate a variable-length parallel generation framework.
Figure 1: Overall framework of the proposed method. The feature extractor transforms the source audio into input features. These features are subsequently disentangled and reconditioned by the style autoencoder. The unit aligner is responsible for providing unit-level context and performing duration modeling. In addition, the hierarchical style encoder encodes features at both the unit and frame levels. Mel-spectrograms are subsequently produced by the generator. In this figure, ”DP” represents the duration predictor, $\mathcal{LR}$ denotes the length regulator, while ”Q”, ”K” and ”V” represent the query, key and value of the cross-attention in the unit aligner, respectively. $\otimes$ denotes the concatenate operation. $w_{src}$ represents the source style vector and $w_{tgt}$ represents the target style vector. The style autoencoder disentangles the source style from the features and applies the target style, while the hierarchical stylize encoder and generator take the target style as a condition.
III Proposed Method
We propose DurFlex-EVC, which is flexible in duration and parallel in generation. The model consists of the following components:
  • Feature extractor that transforms raw audio waveforms into acoustic features.
  • Style autoencoder, facilitating the disentanglement of content and style.
  • Unit aligner that handles context transformation and precise duration modeling.
  • Hierarchical stylize encoder that operates at both unit and frame levels to capture stylization.
  • Diffusion-based generator that produces high-quality Mel-spectrograms.
Fig. 1 illustrates the comprehensive framework of DurFlex-EVC, with detailed explanations provided in the following subsections.
III-A Overview
Fig. 1a shows the overall structure of the model. First, the waveform is transformed by the feature extractor to be used as input to the model. Features such as Mel-spectrograms or SSL representations can be used as input. In our method, we adopted HuBERT [29] as a feature extractor and used the output of the last layer as input to the model. The style of the input feature is disentangled and is adapted by the style autoencoder. The unit aligner then aggregates contextual information at the frame level through a cross-attention module. This representation is compressed to the unit level, feeding into the hierarchical stylize encoder. Features are stylized at the unit-level, and then stylized after frame-level extension. The diffusion-based generator produces a Mel-spectrogram from the output of the hierarchical stylize encoder and a style vector. This Mel-spectrogram is then converted into a raw waveform by a pre-trained vocoder.
III-B Style Autoencoder
The feature extractor generates features that include both content and style. In this context, content refers to the linguistic information of the speech, while style encompasses all other aspects, including emotional expression and speaker-specific properties. To model styles, we distinguish between speaker style and emotional style. The style autoencoder is designed to decompose the source emotion style. Fig. 1b shows the structure of the style autoencoder, which consists of two primary components: the de-stylize transformer and the stylize transformer. The de-stylize transformer causes the style to be decomposed from the feature, while the stylize transformer applies the style. The de-stylize transformer is shown in Fig. 1f and the stylize transformer in Fig. 1g. These components have been designed using advanced normalization techniques.
Firstly, layer normalization (LN) serves as a fundamental technique, mathematically expressed as follows:
$\text{LN}(z)=\frac{z-\mu}{\sigma},$ (1)
where $z$ represents the input vector to be normalized, $\mu$ is the mean of the vector, and $\sigma$ is its standard deviation.
The stylize transformer employs conditional layer normalization (CLN) [58] based on LN to effectively adapt styles. CLN is defined as follows:
$\text{CLN}(z,w)=\gamma(w)\times\text{LN}(z)+\beta(w),$ (2)
where $\gamma(w)$ and $\beta(w)$ are adaptive parameters representing the gain and bias for the style vector $w$.
In contrast, the de-stylize transformer employs mix-style layer normalization (MixLN) [59], an modification of CLN, to disentanlge style-independent features. MixLN introduces perturbations in the input directed towards the style vector, inhibiting the model’s tendency to learn style-specific features. This perturbation is executed by blending the original style vector with batch-level shuffled style vector.
$\gamma_{mix}(w)=\lambda\gamma(w)+(1-\lambda)\gamma(\tilde{w}),$ (3)
$\beta_{mix}(w)=\lambda\beta(w)+(1-\lambda)\beta(\tilde{w}),$ (4)
where $w$ and $\tilde{w}$ denote the original and shuffled style vector, respectively. The $\lambda$ is responsible for balancing the original style with the shuffle style and follows a beta distribution, $\text{Beta}(\alpha,\alpha)$ with $\alpha\in(0,\infty)$. This parameter resides in a $B$-dimensional real number space, denoted as $\mathbb{R}^{B}$, where $B$ represents the batch size, the amount of data processed simultaneously by the model. Therefore, MixLN is defined as follows:
$\text{MixLN}(z,w)=\gamma_{mix}(w)\times\text{LN}(z)+\beta_{mix}(w).$ (5)
A style autoencoder is composed of $N$ de-stylize transformers and $N$ stylize transformers. We constructed the source style vector $w_{src}$ as the sum of the speaker vector $s_{src}$ and the emotion vector $e_{src}$, and the target style vector $w_{tgt}$ as the sum of the speaker vector $s_{src}$ and the emotion vector $e_{tgt}$.
$w_{src}$ $=s_{src}+e_{src},$ (6)
$w_{tgt}$ $=s_{src}+e_{tgt}.$ (7)
The speaker vector $s_{*}$ and the emotion vector $e_{*}$ are obtained from the embedding look-up table. The de-stylize transformer uses $w_{src}$ to disentangle the source emotion style from input feature, whereas the stylize transformer utilizes $w_{tgt}$ to apply the target emotion style.
(a) unit-level pooling
(b) frame-level scaling
Figure 2: Unit-level pooling and frame-level scaling. (a) Latent is pooled on average based on unit durations, and (b) Latent is expanded by being duplicated a number of times corresponding to the duration count.
III-C Unit Aligner
The unit aligner is introduced to model the semantic context in which frame-level features are compressed to the unit level, which works similar to an information bottleneck. Fig. 1d shows the structure of the unit aligner. We combined learnable embeddings with cross-attention to derive attention weights for specific embeddings a feature focuses on, which is used for duration modeling. The attention weights are trained to predict the unit through the index of the largest value, and the duration is modeled as the consecutive number of units obtained. We use the output of the style autoencoder as a query ($Q$) and introduce learnable embeddings $e_{unit}$ as keys ($K$) and values ($V$) to cross-attention based on [60]. The attention weights $\mathcal{A}_{unit}$ are computed as follows:
$\mathcal{A}_{unit}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right),$ (8)
where $d$ is the dimension of $Q$ and $K$. Subsequently, these weights $\mathcal{A}_{unit}$ are integrated with the value matrix $V$ to produce the attention output $z_{attn}$.
$z_{attn}=\mathcal{A}_{unit}\cdot V.$ (9)
Fig. 1e shows the process of learning attention weights to predict units and the modeling of the duration of the predicted units. We introduce an additional loss term to guide the attention module to learn semantic information. This approach implies a direct classification task, correlating the attention weights $\mathcal{A}_{unit}$ with the target unit sequence $y$.
$\mathcal{L}_{unit}=-\frac{1}{L}\sum_{i=1}^{L}\sum_{j=1}^{C}y_{i,j}\log(\mathcal{A}_{unit}^{i,j}),$ (10)
where $L$ is the length of the unit sequence and $C$ is the number of unit classes, $\mathcal{A}_{unit}^{i,j}$ represents the predicted probability of the $i$-th element for class $j$ from the attention weights $\mathcal{A}_{unit}$, and $y_{i,j}$ is the one-hot encoded class label for the $i$-th unit for class $j$. We adopted the HuBERT unit as our target unit. This is an intended design feature that ensures that the context reflected in the unit aligner is consistent with the target style, rather than based on the context of the input speech. During emotional conversion, the style autoencoder removes the source emotion from the input feature and applies the target emotion. The unit aligner then forms a unit-level context for the target emotional style.
III-D Duration Modeling
The unit sequence can be predicted by identifying the focus of the attention module.
$\hat{y}_{i}=\arg\max(\mathcal{A}_{unit}^{i}),$ (11)
where $\hat{y}^{i}$ is the $i$-th predicted unit and $\mathcal{A}_{unit}^{i}$ is the $i$-th frame of the attention weights. To extract a distinct sequence of units and their consecutive counts, a deduplication operation is applied, represented as:
$\hat{y}_{uniq},n_{count}=\text{dedup}(\hat{y}),$ (12)
where $\hat{y}_{uniq}$ and $n_{count}$ denote unique units and their consecutive counts, respectively. For example, given an input sequence $\hat{y}=[4,4,2,2,2,2,1,1]$, the deduplication operation results in $\hat{y}_{uniq}=[4,2,1]$ and $n_{count}=[2,4,2]$. This implies two units with index $4$, followed by four units with index $2$ and two units with index $1$. We train the duration predictor with $n_{count}$ as the target duration. This duration is used to perform unit-level pooling to semantically bundle the output of the unit aligner. Fig. 2a explains unit-level pooling. The output of the unit aligner, $z_{attn}$, is averaged based on the duration of the unit and results in downsampling the sequence length for alignment at the unit level.
$z_{u}=\text{unit-level-pooling}(z_{attn},n_{count}),$ (13)
where $z_{u}$ is the latent downsampled at the unit level. For example, given that $z_{attn}=[0.2,0.2,0.1,0.4,0.5,0.2,0.3,0.5]$ and $n_{count}=[2,4,2]$, the result of unit-level pooling is $z_{u}=[0.2,0.3,0.4]$.
III-E Hierarchical Stylize Encoder
The hierarchical stylize encoder [61] functions at two levels: unit level and frame level. It consists of two components: the unit-level stylize transformer (UST) and the frame-level stylize transformer (FST). UST processes $z_{u}$ into a $z_{us}$, denoted as $z_{us}=\text{UST}(z_{u},w_{tgt})$, focusing on unit-specific features. This refined variable $z_{u}$ is scaled to the frame level through a length regulator $\mathcal{LR}$, depicted in Fig. 2b.
$z_{f}=\mathcal{LR}(z_{us},n_{count}),$ (14)
where $z_{f}$ represents the latent variable at the frame level. For example, if $z_{us}=[0.1,0.2,0.5]$ and $n_{count}=[2,5,1]$, then $z_{f}$ becomes $[0.1,0.1,0.2,0.2,0.2,0.2,0.2,0.5]$. The FST further refines the frame-level features $z_{f}$ to $z_{fs}$, expressed as $z_{fs}=\text{FST}(z_{f},w_{tgt})$. This final output $z_{fs}$ is subsequently used as input for the Mel-spectrogram generator.
The duration predictor takes $z_{us}$ and $w_{tgt}$ as input and is trained to predict the unit-level duration $n_{count}$. For emotion-based duration dynamics, we introduce the flow-based stochastic duration predictor proposed in [62] to introduce duration uncertainty. The duration predictor training objective $\mathcal{L}_{dur}$ follows a negative variational lower bound.
III-F Diffusion-Based Mel-Spectrogram Generator
We use a diffusion framework based on stochastic differential equations (SDE) to generate high-quality speech with expressive emotions. The diffusion-based model gradually transforms the Mel-spectrogram into Gaussian noise in a forward process and generates samples from the noise in a reverse process. We adopt the standard normal distribution as the prior distribution, as in [25]. The model is trained to minimize the mean square error (MSE) loss $\mathcal{L}_{diff}$ between the ground truth noise and the estimated noise. For score estimation, our model incorporates a network denoted by $s_{\theta}$ based on the U-net architecture with linear attention used in Grad-TTS [63].
III-G Training Objective
Consequently, the model is trained using the following loss function:
$\mathcal{L}_{total}=\lambda_{diff}\mathcal{L}_{diff}+\lambda_{unit}\mathcal{L}_{unit}+\lambda_{dur}\mathcal{L}_{dur},$ (15)
where $\lambda_{diff}$, $\lambda_{unit}$, and $\lambda_{dur}$ are the loss weights, which we set to 1.0, 0.1, and 0.1, respectively.
III-H Emotion Voice Conversion Process
The process of converting the emotion in the input speech to the target emotion is as follows.
  1. 1.
    The input waveform is converted into input features by the feature extractor.
  2. 2.
    The features are de-stylized from the source style vector and stylized to the target style vector by the style autoencoder. The source style vector is obtained from the source emotion vector and the speaker vector. The target style vector is obtained from the target emotion vector and the speaker vector. The source style is disentangled by the MixLN of the style autoencoder, and the target style is applied by the CLN.
  3. 3.
    Unit-level features according to the target style are obtained with the unit aligner.
  4. 4.
    The hierarchical stylize encoder adapts the target style to the features at the unit-level and the frame-level.
  5. 5.
    The diffusion-based generator produces a Mel-spectrogram conditioned on the feature and target style vector.
  6. 6.
    The waveform is synthesized by pre-trained vocoder.
IV Experiments
IV-A Experimental Setup
We conducted experiments using the emotional speech dataset (ESD) [^1] https://github.com/HLTSingapore/Emotional-Speech-Data[64], which contains 350 parallel utterances spoken by 10 native Mandarin speakers and 10 English speakers with 5 emotional states (neutral, happy, angry, sad, and surprise). Following the data partitioning guidelines provided by ESD, we constructed the training set with 300 samples per emotion per speaker, for a total of 15,000 samples. The validation set included 20 samples for each emotion per speaker, totaling 1,000 samples, and the test set comprised 30 samples for each emotion per speaker, totaling 1,500 samples. We sampled audio at 16,000 Hz and transformed it to an 80-bin Mel-spectrogram using a short-time Fourier transform (STFT) with a window length of 1,024 and a hop size of 256.
The experiments included transformations between all possible emotional states, not just limited from neutral to other states. This approach was designed to cover all possible emotional state conversions, ensuring a comprehensive assessment of the model’s performance. For subjective evaluation, 10 sentences were randomly selected for each of the 5 emotions. These sentences were then adapted to reflect each of the four other emotional states, resulting in a total of 200 samples ($10\times 5\times 4=200$). For objective evaluation, each of the 1500 test samples was transformed into the four other emotional states, resulting in a total of 6000 samples ($1500\times 4=6000$). This setup provided a thorough assessment of the model’s performance across all possible emotional state conversions.
IV-B Implementation Details
In our experimental setup, we configured both the de-stylize and stylize transformers with specific parameters: the hidden dimension, kernel size, number of heads, FFN kernel size, and feed forward network (FFN) hidden size were set to 256, 5, 2, 9, and 1024, respectively. The $\alpha$ parameter of the Beta distribution for MixLN was fixed at 0.1. All transformers used in our model were organized into $N$ layers, with $N$ established at 4. The unit aligner featured multi-head attention with 16 heads. We set $T=1$, $\beta_{t}=\beta_{0}+(\beta_{1}-\beta_{0})t$, $\beta_{0}=0.05$, and $\beta_{1}=20$ as noise schedules. The U-Net in our model was set to downsample four times and had a hidden dimension of 128. We set the inference timestep to 100. The duration predictor, which comprises residual blocks using dilated and depth-separable convolution, was structured in four layers. To address the resolution disparity between the HuBERT unit and the Mel-spectrogram, we expanded the hidden representation with a length regulator and employed linear interpolation for upsampling. In training the generator, we utilized random segments, setting the segment size to 32 frames of the Mel-spectrogram. The AdamW optimizer was used, with a learning rate of $1\times 10^{-4}$. We set the batch size to 16 and the training steps to 500K. We trained the vocoder using the official BigVGAN [^2] https://github.com/NVIDIA/BigVGAN[65] implementation, incorporating LibriTTS [66], VCTK  [^3] https://datashare.ed.ac.uk/handle/10283/2651, and ESD datasets. All comparison models were trained using a single NVIDIA RTX A6000 GPU. For broader accessibility, the code [^4] https://github.com/hs-oh-prml/DurFlexEVC and a demo [^5] https://prml-lab-speech-team.github.io/durflex/ of our proposed method are available online.
IV-C Evaluation
IV-C1 Subjective Metrics
We conducted subjective evaluations using Amazon Mechanical Turk (mTurk). Our analysis included the mean opinion score (MOS) for naturalness (nMOS) and speaker similarity (sMOS), using a 9-point scale, ranging from 1 to 5, with increments of 0.5 units. The results are presented with a confidence interval (CI) of 95%. Furthermore, we use the emotion mean opinion classification (eMOC) as suggested in [42].
TABLE I: Performance of the pre-trained evaluator model on test sets
Model UTMOS PER CER WER ECA EECS SECS
GT 3.60 11.64 3.06 12.09 96.33 0.93 0.81
GT (vocoded) 3.58 11.73 3.14 12.45 94.13 0.91 0.81
(a) Speaker
(b) Emotion
Figure 3: Visualize t-SNE of emotion2vec features for speaker and emotion.
(a) StarGAN-EVC
(b) Seq2seq-EVC
(c) Emovox
(d) Mixed-Emotions
(e) Textless-EVC
(f) DurFlex-EVC
Figure 4: Comparison of the SECS scores of the comparison models for all combinations of emotion conversion.
(a) StarGAN-EVC
(b) Seq2seq-EVC
(c) Emovox
(d) Mixed-Emotions
(e) Textless-EVC
(f) DurFlex-EVC
Figure 5: Comparison of the EECS scores of the comparison models for all combinations of emotion conversion.
IV-C2 Objective Metrics
For our objective evaluation, we incorporate a range of metrics: predicted mean opinion score, phoneme error rate (PER), character error rate (CER), word error rate (WER), emotion classification accuracy (ECA), and speaker embedding cosine similarity (SECS). The predicted MOS was assessed using UTMOS [67] [^6] https://github.com/tarepan/SpeechMOS. For the PER calculation, we used a wav2vec2.0-based phoneme recognition model from Hugging Face [68]. CER and WER were determined using Whisper [^7] https://github.com/openai/whisper[69]. In assessing SECS, we extracted speaker embeddings from both target and generated audio using Resemblyzer [^8] https://github.com/resemble-ai/Resemblyzer, subsequently computing their cosine similarity. This similarity measure ranges from -1 to 1, where higher values denote greater similarity. We evaluated the similarity for samples that shared the same speaker and emotion and then averaged these across all speakers. The objective evaluation of emotions in the generated speech was conducted using a pre-trained speech emotion recognition (SER) model. To measure SER accuracy, we employed emotion2vec [70]. We used emotion2vec+ base [^9] https://github.com/ddlBoJack/emotion2vec, a pre-trained model that supports nine classes, and only used the five sentiment classes in the ESD dataset for evaluation. We propose the emotion embedding cosine similarity (EECS) to evaluate the emotion of synthesized speech. The EECS is obtained by computing the cosine similarity of the emotion embedding between the synthesized audio and arbitrary reference audio with the target emotion. The emotion embedding was obtained using emotion2vec. Fig. 3 is a visualization of the features in emotion2vec, which shows that it encodes emotions independently of the speaker. We also evaluated the root mean square error (RMSE) for pitch and energy and calculated the difference of duration (DDUR) to assess prosody. The pitch was extracted using parselmouth [^10] https://parselmouth.readthedocs.io/en/stable/ in Hz, and the energy was obtained as the L2-norm of the absolute value of the linear spectrogram. DDUR was obtained in the same way as in [22]. Table I presents the results of each pre-trained baseline model on the test set, comprising 1500 samples, for both ground-truth and vocoded samples.
IV-D Comparison Models
To benchmark the efficacy of our proposed method, we trained and compared it against several existing models.
  • StarGAN-EVC [^11] https://github.com/glam-imperial/EmotionalConversionStarGAN[15]: This adversarial network model specializes in speech emotion conversion. Its GAN-based architecture supports parallel generation, distinguishing it in this domain.
  • Seq2seq-EVC [^12] https://github.com/KunZhou9646/seq2seq-EVC[21]: Employing a sequence-to-sequence (seq2seq) framework, this model adopts a two-stage strategy utilizing the TTS model. A notable feature of Seq2seq-EVC is its ability to jointly model duration and pitch.
  • Emovox [^13] https://github.com/KunZhou9646/Emovox[22]: Similar to Seq2seq-EVC, Emovox is based on a seq2seq structure. Its uniqueness lies in its focus on modulating emotional intensity. Emovox incorporates a ranking function to effectively model this intensity dimension.
  • Mixed-Emotions [^14] https://github.com/KunZhou9646/Mixed_Emotions[71]: Operating on a seq2seq framework similar to Emovox, this model is designed to express mixed emotions. It shares a controllable emotion intensity feature with Emovox.
  • Textless-EVC [^15] https://github.com/facebookresearch/fairseq/tree/main/examples/emotion_conversion[42]: This model approaches speech synthesis by deconstructing the speech signal into discrete learned representations. These include speech content units, prosodic features, speaker identity, and emotions. Each element is modified to align with the target emotion before being synthesized back into speech.
  • DurFlex-EVC: Our proposed model includes a unit aligner, style autoencoder, stochastic duration predictor, hierarchical stylize encoder, and a diffusion-based generator. This model stands out with its comprehensive and integrated approach to emotional speech synthesis.
All comparison models were trained using the official implementation. We used the same vocoder to generate the waveforms except for Textless-EVC, which generates the waveform directly, and adjusted the hyperparameters to match the vocoder settings.
TABLE II: Results of subjective evaluations each comparison model
Model nMOS sMOS eMOC
GT 3.72 ($\pm$0.03) 3.95 ($\pm$0.06) 82.98
GT (vocoded) 3.70 ($\pm$0.05) 3.58 ($\pm$0.11) 82.98
StarGAN-EVC 3.59 ($\pm$0.06) 3.36 ($\pm$0.12) 37.84
Seq2seq-EVC 3.43 ($\pm$0.07) 3.09 ($\pm$0.13) 48.65
Emovox 3.50 ($\pm$0.06) 3.10 ($\pm$0.13) 51.35
Mixed Emotion 3.50 ($\pm$0.07) 3.27 ($\pm$0.12) 62.16
Textless-EVC 3.61 ($\pm$0.05) 3.39 ($\pm$0.11) 56.76
DurFlex-EVC 3.70 ($\pm$0.05) 3.63 ($\pm$0.10) 72.97
TABLE III: Results of objective evaluations for each comparison model
Model UTMOS PER CER WER ECA EECS SECS
StarGAN-EVC 1.47 70.83 44.49 67.71 39.5 0.34 0.61
Seq2seq-EVC 1.54 37.29 21.68 36.87 40.0 0.39 0.63
Emovox 2.05 29.25 17.18 31.37 49.33 0.48 0.68
Mixed Emotion 2.02 29.86 18.21 33.09 57.75 0.55 0.67
Textless-EVC 2.37 22.88 12.49 23.98 56.18 0.58 0.68
DurFlex-EVC 3.39 17.31 8.26 20.75 88.64 0.85 0.75
V Result
This section contains the results and discussion of the extensive experiments. We compared our proposed model with previous EVC models to evaluate its quality. We then conducted experiments to demonstrate the effectiveness of the components of the model. Furthermore, we conducted extended experiments on unseen speaker scenarios.
V-A Comparison of evaluation results for baseline models
To evaluate the performance of the proposed model, we conducted both objective and subjective evaluations to compare it with the baseline models. Table II shows the results of subjective evaluation. Table III shows the objective evaluation results. As the nMOS results show, the naturalness of the speech generated by our method outperforms other models. In terms of speaker similarity, our approach scored the highest on both sMOS and SECS. Fig. 4 shows the resulting SECS for all combinations of emotional conversions. This shows that the proposed model is more robust in terms of speaker similarity than the comparison models for all combinations of transformations. This means that our model is more robust in terms of speaker similarity than the comparison models for all combinations of transformations. Furthermore, the results of eMOC, ECA and EECS demonstrate that our method performs better in terms of perceptual quality as well as objective metrics. Fig. 5 shows the resulting EECS for all combinations of emotion transformations. This indicates that the proposed model synthesized speech with a higher emotional similarity than the comparison models for all combinations of transformations. The ASR evaluation also shows that our method achieves lower values in PER, CER, and WER compared to other models. This emphasizes the ability of our method to synthesize precisely pronounced speech. For a quantitative assessment of prosody, we also compared pitch and energy duration. Table LABEL:prosody shows the evaluation results for each emotion. We found that the proposed model scored better than the other models in all prosody evaluations.
TABLE IV: Comparison of prosody features (pitch, energy, duration)
Pitch
Model Neutral Angry Happy Sad Surprise Avg.
StarGAN-EVC 54.96 46.5 59.14 53.51 68.69 56.56
Seq2Seq-EVC 62.03 56.6 70.78 55.21 65.09 61.94
Emovox 53.35 49.35 57.18 51.42 52.96 52.85
Mixed-Emotion 55.0 49.24 56.92 51.63 55.01 53.56
Textless-EVC 57.15 47.85 51.09 54.86 48.72 51.93
DurFlex-EVC 49.43 45.79 50.53 47.46 54.39 49.52
Energy
Model Neutral Angry Happy Sad Surprise Avg.
StarGAN-EVC 21.00 19.80 20.17 20.92 19.54 20.28
Seq2Seq-EVC 25.46 24.54 24.30 24.67 24.32 24.66
Emovox 25.22 24.64 24.02 24.73 24.01 24.53
Mixed-Emotion 25.46 24.20 23.74 24.40 23.56 24.27
Textless-EVC 13.44 13.52 14.32 13.96 13.88 13.82
DurFlex-EVC 12.25 12.31 12.64 13.16 12.63 12.60
Duration
Model Neutral Angry Happy Sad Surprise Avg.
StarGAN-EVC 0.31 0.34 0.26 0.26 0.28 0.29
Seq2Seq-EVC 0.21 0.22 0.27 0.22 0.23 0.23
Emovox 0.22 0.23 0.23 0.21 0.24 0.23
Mixed-Emotions 0.22 0.22 0.29 0.25 0.30 0.26
Textless-EVC 0.30 0.36 0.34 0.28 0.33 0.32
DurFlex-EVC 0.20 0.21 0.24 0.23 0.23 0.22
Figure 6: Cross-entropy change across layers for emotion classification. Models with the SAE show increasing cross-entropy through de-stylize transformers and decreasing cross-entropy through stylize transformers. Models without the SAE show consistently high cross-entropy.
TABLE V: Results of ablation studies and additional experiments
Model UTMOS PER CER WER ECA EECS SECS
DurFlex-EVC 3.39 17.31 8.26 20.75 88.64 0.85 0.75
w/o SAE 3.34 18.28 9.37 22.64 87.64 0.83 0.72
w/o UA 3.55 12.31 3.55 13.04 24.39 0.31 0.66
w/o HSE 3.27 20.00 9.32 22.65 68.95 0.65 0.69
w/ DDP 3.39 17.60 8.59 21.56 87.52 0.83 0.73
w/ FFT 3.03 17.32 7.42 19.41 54.68 0.57 0.73
w/ adv 3.38 18.47 8.33 21.18 87.53 0.84 0.71
w/ unit2mel 3.30 20.18 9.32 22.65 89.11 0.85 0.69
w/ unit2wav 1.26 18.86 7.69 18.56 20.23 0.29 0.51
V-B Experiments for Analyzing Model Architectures
We conducted an analysis of the model design and additional experiments. Table V includes the results of the ablation study and additional experiments.
V-B1 Effectiveness of Style Autoencoder
We conducted an experiment to verify the effectiveness of the style autoencoder (SAE). w/o SAE is a model that stacks a standard feed forward transformer block with a full layer of SAE, without purtubation and conditioning for emotion style. In Table V, w/o SAE shows the results of the ablation study on it. The overall performance degradation observed in models without SAE indicates the importance of disentangling content and style from input features. Fig. 6 shows the change in cross entropy for styles as input features pass through SAE. To obtain cross entropy, we trained a classifier for emotion style on features extracted from all layers of SAE. The black line is the result for the model with SAE and the orange line is the result for the model without SAE. The w/o SAE model shows high cross entropy across the layers, while the model with SAE shows increasing cross entropy as it passes through the de-stylize transformer layers and decreasing cross entropy as it passes through the stylize transformer layers. This means that the style is disentangled and conditioned from the feature by SAE.
We also experimented with adversarial training strategies to remove source styles from input features. The results for this are shown in Table V as w/ adv, and show a performance degradation on all metrics except EECS. This indicates that the proposed method is more stable for model learning than traditional adversarial learning.
TABLE VI: JSD of unit distributions by emotion
Neutral Angry Happy Sad Surprise
Textless-EVC 0.21 0.20 0.22 0.22 0.21
DurFlex-EVC 0.11 0.11 0.12 0.12 0.11
V-B2 Effectiveness of Unit Aligner
The unit aligner (UA) was introduced to model stylized contexts. Table V w/o UA shows the results of the ablation experiment for this. The results showed that w/o UA performed better in terms of voice quality and pronunciation, but not in terms of emotion conversion. This indicates that there remains a lot of information about the source style in the feature and that the UA functions as a bottleneck and has a significant impact on style control. We measured the Jensen-Shannon divergence (JSD) to validate that the predicted units in UA reflect the target emotion. Table VI shows the JSD results of the units for each emotion in Textless-EVC and DurFlex-EVC. DurFlex-EVC achieved a better JSD than Textless-EVC for all emotions, indicating that it provides an appropriate representation of the context.
V-B3 Effectiveness of Hierarchical Stylize Encoder
The output of the unit aligner is a representation in units, which is expanded to frame-level by a duration predictor and length regulator. For frame-level stylization in this representation, we introduce an heirarchical stylize encoder (HSE), and the results of the ablation study are shown in Table V w/o HSE. The evaluation results showed an overall performance degradation with w/o HSE. This is inferred to be due to HSE reducing the burden on the Mel-spectrogram generator, resulting in improved generation quality.
(a) Angry
(b) Happy
(c) Sad
(d) Surprise
Figure 7: Histogram of duration by emotion based on the structure of the duration predictor.
TABLE VII: JSD for unit durations by emotion
Model Neutral Angry Happy Sad Surprise
w/ DDP (w/o dropout) 0.037 0.021 0.021 0.064 0.021
w/ DDP 0.037 0.021 0.022 0.063 0.022
w/ SDP 0.027 0.015 0.017 0.050 0.021
V-B4 Comparison for Duration Predictors
We introduced a stochastic duration predictor (SDP) for duration modeling to represent the diversity of emotions. To evaluate, we compared our experiments with widely used deterministic duration predictors (DDP). The DDP follows the structure described in FastSpeech [43], which includes two convolutional layers, ReLU activation, layer normalization, and dropout. Tables V, w/ DDP show the evaluation results of the model using DDP. The evaluation results show that the model using SDP is better than the model using DDP. We compared the JSD of unit duration for each emotion for each duration predictor. Table VII shows the JSD results for unit duration for models with stochastic duration predictors, w/ SDP, and deterministic duration predictors, w/ DDP.
In our experiments, we found that DDP does not always output the same length when generating the same sentence with the same conditions. We discovered that it was caused by a dropout within the DDP. We experimented with repeatedly generating the same speech, and Fig. 7 shows a histogram of the duration of the speech for each emotion. Red bars represent DDP, black bars represent SDP, and ’x’ represents the results for DDP with dropout removed. Table VII, w/ DDP (w/o dropout), shows the results for a fully deterministic duration predictor without dropout, which gave almost similar results to the version with dropout. SDP showed better JSD than w/ DDP and w/ DDP (w/o dropout), indicating that it is suitable for modeling emotional duration distributions.
(a) DurFlex-EVC w/ FFT
(b) DurFlex-EVC w/ diffusion
Figure 8: Pitch track of the same speech converted multiple times for each emotion.
TABLE VIII: JSD for pitch by emotion
Model Neutral Angry Happy Sad Surprise
w/ FFT 0.21 0.13 0.15 0.17 0.24
w/ diffusion 0.19 0.11 0.16 0.15 0.25
V-B5 Comparison for Mel-spectrogram Generator Structures
Diffusion-based generators have been shown to produce high-quality and diverse results in a wide range of domains. We introduced a diffusion-based structure to generate more expressive speech for each emotion. We conducted experiments to compare our decoder with a feed-forward transformer (FFT) based structure, which is widely used in conventional speech synthesis studies for parallel generation. Table V w/ FFT shows the objective evaluation results of the model using the FFT-based decoder. The w/ FFT scored better in CER and WER for pronunciation, but lower on UTMOS for quality, and worse on ECA and EECS for emotion expression.
To verify the expressiveness of each emotion, we computed the JSD over pitch for each emotion. In Table VIII, w/ FFT is the model using an FFT decoder and w/ diffusion is the model using a diffusion-based structure. w/ diffusion shows better JSD on neutral, angry, sad, while w/ FFT shows better JSD in happy and surprise. Although w/ FFT outscores w/ diffusion in Happy and Surprise, it does not mean that FFT is more expressive. Fig. 8 shows the pitch of the speech generated by multiple iterations of the same sentence. Models using FFT produce consistent pitch tracks that remain stable over repeated generations, while models using diffusion produce more dynamic results. It can be interpreted that the diffusion-based structure models a higher expressiveness, resulting in a higher JSD than the FFT-based structure, which models the average over dynamic emotions (happy, sad).
TABLE IX: Comparison of results based on input features
Model UTMOS PER CER WER ECA EECS SECS
w/ Mel-spec. 3.29 25.28 15.70 31.31 88.10 0.84 0.72
w/ linear-spec. 3.29 25.98 16.29 31.81 89.88 0.85 0.68
w/ wav2vec 2.0 3.34 30.55 21.46 38.44 91.03 0.86 0.66
w/ wavLM 3.36 23.44 12.07 26.22 92.59 0.87 0.67
w/ HuBERT 3.39 17.31 8.26 20.75 88.64 0.85 0.75
TABLE X: Comparison of BLEU score and UER based on input features
Model BLEU UER
w/ Mel-spec. 15.23 62.62
w/ linear-spec. 15.66 60.93
w/ wav2vec 2.0 11.96 62.67
w/ wavLM 25.51 46.43
w/ HuBERT 38.59 38.04
V-B6 Comparison for Input Features
The proposed model takes HuBERT features as input and outputs a Mel-spectrogram. The reason for outputting Mel-spectrogram is for compatibility with pre-trained vocoders. To explain the reasoning behind our choice of HuBERT features, we present a comparison of the input features. We compared Mel-spectrogram and linear spectrogram as features using conventional signal processing, and wav2vec 2.0 [27], wavLM [32], and HuBERT [29] as SSL features that have recently been used as linguistic representations [35, 33]. Table IX shows the results of the experiments by input feature. The results show that models using Mel-spectrogram and linear-spectrogram have lower UTMOS than SSL feature models. Models using wavLM performed better overall than those using wav2vec 2.0. The model using HuBERT outperformed the others in all metrics except ECA and EECS. In particular, the model using HuBERT shows a significant improvement in pronunciation, which is interpreted as advantageous for linguistic learning over other features because the target units used for training are obtained from the clustering of HuBERT features. To compare the linguistic modeling ability of each input feature, we calculated the BLEU score [72] and the unit error rate (UER) of the prediction unit. Table X shows the BLEU score and UER of the predicted units for each feature. We found that the prediction of units was correlated with the pronunciation accuracy of the synthesized speech and that this was due to the fact that HuBERT predicted units more accurately than the other features.
We also experimented with a model that takes unit input and generates speech. In Table V, w/ unit2mel is a model that receives unit input and generates Mel-spectrogram. The w/ unit2wav is a model that predicts units and generates waveforms. The w/ unit2mel shows a poorer overall performance, except for ECA and EECS. The w/ unit2wav shows significantly worse metrics across the board, except for CER and WER. We interpret this to mean that the speech unit has enough information about pronunciation but not enough other speech information to generate a waveform. To overcome this, Textless-EVC uses additional information such as pitch and timbre to generate the waveform.
Figure 9: Style embedding modeling for unseen speaker setting.
Figure 10: Pitch track of unseen speaker speech converted to each emotion.
TABLE XI: Evaluation results for seen and unseen speakers
Model UTMOS PER CER WER ECA EECS SECS
GT (Seen) 3.60 11.64 3.06 12.09 89.46 0.76 0.81
GT (Unseen) 4.03 10.07 0.67 1.39 - - 0.84
Seen 3.44 16.72 7.75 20.16 82.92 0.75 0.66
Unseen 3.53 18.83 7.35 12.18 75.85 0.72 0.60
V-C Unseen Speaker Emotion Conversion
We extended our experiments to apply our proposed model to an unseen speaker scenario. To make it possible, we modified the model structure to allow speaker information to encode speaker embedding from reference audio instead of speaker IDs. We adopt the style encoder structure from Meta-StyleSpeech [73] as the speaker encoder. We added a gradient reversal layer (GRL) [74] and a linear layer to prevent the speaker encoder from learning information about emotion. The linear layer performs the emotion classification task, where the losses are reversed by the GRL to prevent the speaker encoder from learning about emotion. Fig. 9 shows the speaker encoder and emotion embedding designed in this way to model style embedding. We set the weight for adversarial losses due to GRL to 0.001. We trained our model on the ESD dataset. For the unseen speaker test, we composed the testset by randomly selecting five sentences for each speaker from the VCTK dataset that were not used for training. We set the emotion of the test set to neutral and converted all other emotions.
Table XI shows the evaluation results for seen and unseen speakers for the modified model. The modified model scored better UTMOS, PER, CER, and WER than the original version, while performing weaker on ECA, EECS, and SECS. We guess that the synthesis quality and pronunciation are better due to the additional information encoded from the reference audio that helps with speech synthesis. However, the lack of style disentanglement leads to decreased performance on emotion and speaker-related metrics. The results of the unseen speaker show that the modified structure allows for the emotion conversion of a new speaker without losing quality. Fig. 10 shows the pitch tracks of the transformed samples of the unseen speaker for each emotion, showing distinct differences for each emotion. We found that the speaker similarity of unseen speakers was poor compared to seen speakers. Table XI shows the speaker similarity for each emotion. We observe a decrease in the speaker similarity for all emotions. Fig. 11 shows a t-SNE visualization of speaker embedding for the ESD dataset (green), the VCTK dataset (purple), and the results of the unseen speaker experiment (yellow). The model synthesized speech that was closer to the speaker in the ESD, which was the training set.
Figure 11: Visualization of t-SNEs of speaker embedding for the ESD dataset (green), the VCTK dataset (purple) and unseen speaker test results (yellow).
TABLE XII: SECS for seen & unseen speaker emotional conversions
Dataset Angry Happy Sad Surprise Avg.
Seen 0.67 0.65 0.66 0.65 0.66
Unseen 0.63 0.58 0.60 0.59 0.60
VI Discussion
Our model is influenced by the framework described in [75], which focuses on modeling emotional pronunciation. However, our approach differs substantially; whereas their model directly addresses pronunciation, our focus is on converting SSL features at the unit level. This perspective aligns more closely with the emotional translation mechanisms inherent in Textless-EVC [42]. The primary distinction in our approach lies in the utilization of the cross-attention output as the input for our model, rather than relying on the predicted units.
We designed our model to generate a Mel-spectrogram. This was more efficient than generating the waveform directly, and allowed us to do more experimentation. For example, Textless-EVC, which generates waveforms directly, spent two weeks training, while our model required three days.
VI-A Limitations
In experimental results, diffusion-based model using demonstrated improved results. However, the limitations of diffusion-based structures are their extensive computational demand and time-consuming nature. We anticipate that this challenge will be alleviated by the advent of the recent fast sampling method[76]. Some applications [77] have been made in speech research. The scope of our experiments also encompassed speaker generalization. Although we successfully observed emotional transformations in the voices of unseen speakers, a discernible lack of speaker similarity was apparent, indicating the need for further refinement. Empirically, our observations suggest that models tested on unseen data often reflect the voice distribution of the data used for their pre-training. Expanding the dataset is expected to increase the representational capacity of the generator, and the performance of zero-shot emotion conversion is also expected to improve. In the configuration of our style autoencoder, we utilized MixLN for de-stylize and CLN for stylize. It was observed that the perturbations introduced by MixLN resulted in a compromise that affected the equilibrium between expressiveness and pronunciation accuracy within our model. In addition, the task of effectively separating style from content remains a significant challenge [78, 79], requiring ongoing research and development to address this issue. Although the diffusion model produces high-quality speech, it has limitations because emotion datasets typically have a 16k sampling rate. The audio super-resolution models such as [80] are expected to solve this problem. Although we only experimented with English, it has the potential to be extended by considering a wide range of languages and combining it with speech-to-speech translations [81]. The expression of emotions differs between people, languages, and cultures, and research is needed to reflect these differences.
VI-B Future Works
Fundamentally, the proposed model depends on the performance of the SSL model because it utilizes discrete units. In addition to semantic units such as HuBERT, we also plan to investigate structures that exploit neural audio codecs such as [39, 38]. Recent research in the field of speech synthesis has increasingly focused on controlling emotion intensity. Some works, such as [82] and [83], have adopted the method of modeling emotion intensity using relative attribute ranking functions. Otherwise, some studies, such as [84], have explored the modeling of intensity through interpolation of embeddings. In addition, various approaches have tried to control the intensity of emotions, such as [85]. Although these studies have shown that modeling emotion intensity is possible, precise control of intensity remains a challenge. In our future work, we explore the direction in which emotional intensity can be controlled. We will also investigate emotional voice conversion for cross-language. Furthermore, it is expected to be applied to singing voice synthesis tasks [86] to express emotions.
VII Conclusion
In this work, we introduced DurFlex-EVC, which generates speech of various durations. We leveraged the discrete speech units of HuBERT to model the contents at the unit-level, and achieved duration flexibility by predicting unit duration and extending it to the frame-level. We propose a style autoencoder to disentangle the source style of input features and apply target styles, and a unit aligner to enable emotional context modeling at the unit level. A hierarchical stylize encoder is introduced for stylistic enhancement of features. Further improvements were achieved by including a stochastic duration predictor and a diffusion-based generator. Experiments demonstrated that DurFlex-EVC outperforms existing EVC models in terms of performance. We also extended the models capabilities to scenarios with unseen speakers, and DurFlex-EVC effectively converted emotions while maintaining a high level of speech quality and pronunciation accuracy. We performed a variety of experiments to validate the effectiveness of our proposed model and believe that it can make a significant contribution to the advancement of emotional speech recognition.